# Overthinking Reduction with Decoupled Rewards and Curriculum Data Scheduling

### v1

@@ Abstract
While large reasoning models trained with critic-free reinforcement learning and verifiable rewards (RLVR) represent the state-of-the-art, their practical utility is hampered by “overthinking”, a critical issue where models generate excessively long reasoning paths without any performance benefit. Existing solutions that penalize length often fail, inducing performance degradation due to a fundamental misalignment between trajectory-level rewards and token-level optimization. In this work, we introduce a novel framework, DeCS, built on our theoretical discovery of two previously unaddressed flaws in current length rewards: (1) the erroneous penalization of essential exploratory tokens and (2) the inadvertent rewarding of partial redundancy. Our framework’s innovations include (i) a first-of-its-kind decoupled token-level reward mechanism that surgically distinguishes and penalizes redundant tokens, and (ii) a novel curriculum batch scheduling strategy to master the efficiency-efficacy equilibrium. Experimental results show DeCS can achieve a dramatic reduction in reasoning tokens by over 50% across seven benchmarks while simultaneously maintaining or even improving performance. It demonstrates conclusively that substantial gains in reasoning efficiency can be achieved without compromising a model’s underlying reasoning power.
Figure 1: Left: Two major flaws of prior practice apply sequence-level length reward without control of training data. Negative advantage values penalize correct high entropy tokens from long sequences while positive ones reward redundant tokens from short sequences; Middle: Flaws of length rewards lead to inferior performance and suboptimal efficiency gains on AIME2024 dataset; Right: DeCS improves pass@1 of base models while reducing $\sim 60\%$ token costs compared to the base model across 7 benchmarks. Experimental details are presented in Appendix E.5.
@@ 1 Introduction
Recent large reasoning models (LRM; Guo et al. (2025); OpenAI (2025); Qwen (2025)) trained with critic-free reinforcement learning (RL) algorithms, such as GRPO (Shao et al., 2024), DAPO (Yu et al., 2025), and REINFORCE++ (Hu et al., 2025a), have demonstrated impressive reasoning capabilities through verifiable outcome rewards. A hallmark of such models is their increased propensity to generate high-entropy tokens (e.g., “wait”, “however”, “alternatively”), which serve to bridge logical transitions between reasoning steps (Wang et al., 2025b). While these tokens reflect active reasoning mechanisms that enhance performance, the propagation of trajectory-level rewards to all tokens can inadvertently encourage prolonged generation led by high-entropy tokens even after reaching a correct answer, a phenomenon known as “overthinking” (Ji et al., 2025). To address this inefficiency without sacrificing reasoning quality, recent approaches incorporate a small length penalty into the correctness reward  (Hou et al., 2025; Su & Cardie, 2025; Aggarwal & Welleck, 2025; Zhang et al., 2025d; Kimi et al., 2025; Wu et al., 2025), using critic-free RL frameworks like GRPO to promote concise yet effective reasoning.
@@ 2.2 Efficient Reasoning With Length Penalties
One of the most straightforward methods is to add a length-based reward along with the fundamental correctness reward to encourage shorter yet correct responses (Hou et al., 2025; Su & Cardie, 2025; Aggarwal & Welleck, 2025). Specifically, if adopting a monotonically decreasing length reward function $f(l)=-\gamma l$ accepting the sequence length $l$ as input, the combined reward is defined as:
@@ 3 On the Limitations of Length-Guided Reasoning Optimization
In this section, we formally reveal two significant limitations of current length-reward driven efficiency reasoning under the representative critic-free RLVR algorithm, GRPO, by analyzing the misalignment between the trajectory-level advantage score and the token-level optimization objective for redundant thinking tokens. Through an analysis of logit dynamics, we demonstrate that this misalignment degrades reasoning performance (§3.2) and fails to reduce early redundancies, thereby limiting potential gains in efficiency (§3.3).
@@ 3.3 Insufficient Efficiency
Figure 2: Overview of the DeCS training pipeline. (1) Decoupled Token-level Reward: We finetune a small language model to detect the necessary reasoning prefix (NRP) from other redundancy, which are separately rewarded to penalize overthinking consistently while maintaining the probability for generating necessary reasoning steps. (2) Curriculum Prompt Schedule: The number of easy prompts $q_{\theta,G}$ grows in step with the progressive decline in remaining redundancy.
@@ 4 DeCS
Given the above analysis, we propose DeCS, which contains three main designs to achieve the highest efficacy-efficiency tradeoff. First, to ensure that redundant tokens are penalized deterministically, we train a small module that precisely identifies NRP components for correct trajectories (§4.1). After that, we design a decoupled token-level reward and differentiate the reward scale for necessary and redundant tokens, to ensure enhanced efficiency without compromising performance (§4.2). Based on the conception of NRP, we propose to prevent aggressive penalization on high-entropy tokens following NRP by refactoring the data distribution of a batch according to the current levels of redundancy incrementally (§4.3). Fig. 2 illustrates the overall algorithm.
@@ 4.1 Detection of NRP
Here, $\bigoplus$ denotes the concatenation of reasoning chunks, and the $c^{*}$-th reasoning chunk is the first to entail the correct answer $y^{*}$. The training details are illustrated in Appendix E.6.
@@ 4.2 Decoupled Reward Assignment
$r_{i,j}=\begin{cases}r_{+}\cdot\bm{1}_{{\bm{o}}_{i}\text{ is correct}}\quad&j\leq K_{o_{i}}^{*}\vee o_{j}\notin o_{\rm think}\vee o_{i,j}=\emptyset\\ (r_{0}+\frac{r_{+}-r_{0}}{L_{i}})\cdot\bm{1}_{{\bm{o}}_{i}\text{ is correct}}\quad&j>K_{o_{i}}^{*}\land o_{j}\in o_{\rm think}\end{cases}$ (9)
@@ 4.3 Curriculum Prompt Schedule
where $\kappa_{m}^{0}$ is the ratio of $q_{\theta,G}$ among the current sampled batch and $\beta$ is a hyperparameter to control the learning progress. As trajectories with zero advantages would not provide any learning signal, we follow Yu et al. (2025) to filter prompts whose $G$ rollouts are all incorrect and fill the batch by over-sampling. This curriculum strategy, designed to be bounded and monotonic, enables smooth adjustment in response to the observed NRP ratio, which aligns with the principle of curriculum learning (Bengio et al., 2009). By setting a moderate value $\beta$ with grid search (see Appendix F.1), DeCS can satisfy the condition elucidated in Theorem 1 to maintain unbiased learning of high-entropy tokens throughout the whole training process. This yields stable convergence with no observed training instability or performance degradation, which is reflected in Fig. 6 and 7.
@@ 5.1 Experiment Setups
We use MATH500 (Lightman et al., 2023), AMC23 (AI-MO, 2024), OlympiadBench (He et al., 2024), AIME2024 (Mathematical Association of America, 2025a) and AIME2025 (Mathematical Association of America, 2025b) as in-domain testbeds, GPQA-Diamond (GPQA-D; Rein et al. (2024)) and LiveCodeBench-v6 (LCB; Jain et al. (2025)) as held-out testbeds, covering math, coding, and science tasks with diverse complexity. We choose ThinkPrune (Hou et al., 2025), TLMRE (Arora & Zanette, 2025), AdaptThink (Zhang et al., 2025b), LC-R1 (Cheng et al., 2025) as baselines, and also include GRPO to serve as a performance reference. For fair comparison, we set the temperature as 0.6, top_p as 0.95, and use a maximum token limit of 16384 suggested by Guo et al. (2025). We conduct 128 rollouts for AIME2024, AIME2025, and AMC23, 16 rollouts for OlympiadBench, MATH500 and GPQA-D, and 10 rollouts for LCB to compute pass@1. We also compute the AES score (Luo et al., 2025a) for a comprehensive assessment of efficiency and efficacy. The details of both metrics are presented in Appendix E.4.
@@ 8 Reproducibility Statement
    Datasets: The training set we used is described in §5.1 and evaluation sets we used are described in Appendix E.2.
@@ A.3 Proof of Theorem 2
Table 2: Ablation study with two major components of DeCS on the DS-1.5B base model. “AS” denotes adaptive data sampling and “DR” denotes the decoupled reward mechanism.
@@ Appendix B Related Work
As is introduced in the main text, length reward is one of the most influential strategies to improve efficiency for large reasoning models (LRM). Su & Cardie (2025) proposes to determine the reward scale coefficient $\gamma$ defined in Eq. 4 by comparing the average accuracy of the current step with that of the reference model. Aggarwal & Welleck (2025) proposes L1, which optimizes the LRM to generate a correct reasoning path within a certain context window. Lyu et al. (2025) also proposes to use length penalties to encourage the model to answer problems with different complexity with adaptive token limits. Cheng et al. (2025) uses a specialized model to separate the necessary reasoning part and the redundant tokens, and reinforce the policy to output the response with the shortest NRP length. However, although they have made great efforts to encourage correct yet short reasoning paths, they do not consider the internal logit dynamics when applying sequence-level length rewards with token-level optimization objectives, and thus suffer from performance degradation.
@@
Appendix C Limitation & Future Work

### v2

@@ Abstract
While large reasoning models trained with critic-free reinforcement learning and verifiable rewards (RLVR) represent the state-of-the-art, their practical utility is hampered by “overthinking”, a critical issue where models generate excessively long reasoning paths without any performance benefit. Existing solutions that penalize length often fail, inducing performance degradation due to a fundamental misalignment between trajectory-level rewards and token-level optimization. In this work, we introduce a novel framework, DeCS, built on our theoretical discovery of two previously unaddressed flaws in current length rewards: (1) the erroneous penalization of essential exploratory tokens and (2) the inadvertent rewarding of partial redundancy. Our framework’s innovations include (i) a first-of-its-kind decoupled token-level reward mechanism that surgically distinguishes and penalizes redundant tokens, and (ii) a novel curriculum batch scheduling strategy to master the efficiency-efficacy equilibrium. Experimental results show DeCS can achieve a dramatic reduction in reasoning tokens by over 50% across seven benchmarks while simultaneously maintaining or even improving performance. It demonstrates conclusively that substantial gains in reasoning efficiency can be achieved without compromising a model’s underlying reasoning power. Code is available at https://github.com/pixas/DECS.
Figure 1: Left: Two major flaws of prior practice apply sequence-level length reward without control of training data. Negative advantage values penalize correct high entropy tokens from long sequences while positive ones reward redundant tokens from short sequences; Middle: Flaws of length rewards lead to inferior performance and suboptimal efficiency gains on AIME2024 dataset; Right: DeCS improves pass@1 of base models while reducing $\sim 60\%$ token costs compared to the base model across 7 benchmarks. Experimental details are presented in Appendix G.5.
@@ 1 Introduction
Recent large reasoning models (LRM; Guo et al. (2025); OpenAI (2025); Qwen (2025)) trained with critic-free reinforcement learning (RL) algorithms, such as GRPO (Shao et al., 2024), DAPO (Yu et al., 2025), and REINFORCE++ (Hu et al., 2025a), have demonstrated impressive reasoning capabilities through verifiable outcome rewards. A hallmark of such models is their increased propensity to generate high-entropy tokens (e.g., “wait”, “however”, “alternatively”), which serve to bridge logical transitions between reasoning steps (Wang et al., 2025b). While these tokens reflect active reasoning mechanisms that enhance performance, the propagation of trajectory-level rewards to all tokens can inadvertently encourage prolonged generation led by high-entropy tokens even after reaching a correct answer, a phenomenon known as “overthinking” (Ji et al., 2025). To address this inefficiency without sacrificing reasoning quality, recent approaches incorporate a small length penalty into the correctness reward  (Hou et al., 2025; Su and Cardie, 2025; Aggarwal and Welleck, 2025; Zhang et al., 2025d; Kimi et al., 2025; Wu et al., 2025), using critic-free RL frameworks like GRPO to promote concise yet effective reasoning.
@@ 2.2 Efficient Reasoning With Length Penalties
One of the most straightforward methods is to add a length-based reward along with the fundamental correctness reward to encourage shorter yet correct responses (Hou et al., 2025; Su and Cardie, 2025; Aggarwal and Welleck, 2025). Specifically, if adopting a monotonically decreasing length reward function $f(l)=-\gamma l$ accepting the sequence length $l$ as input, the combined reward is defined as:
@@ 3 On the Limitations of Length-Guided Reasoning Optimization
In this section, we formally reveal two significant limitations of current length-reward driven efficiency reasoning under the representative critic-free RLVR algorithm, GRPO, by analyzing the misalignment between the trajectory-level advantage score and the token-level optimization objective for redundant thinking tokens. Through an analysis of logit dynamics, we demonstrate that this misalignment degrades reasoning performance (§3.2) and fails to reduce early redundancies, thereby limiting potential gains in efficiency (§3.3). The concepts for each involved notation and abbreviation are illustrated in Table 4.
@@ 3.3 Insufficient Efficiency
Figure 2: Overview of the DeCS training pipeline. (1) Decoupled Token-level Reward: We finetune a small language model to detect the necessary reasoning prefix (NRP) from other redundancy, which are separately rewarded to penalize overthinking consistently while maintaining the probability for generating necessary reasoning steps. As the running example “What is 2+3?” shows, the NRP contains the reasoning chunks from the starting token to the first time the model generates the correct answer “5”. After that, any leading redundant token like “Wait” receives negative advantages, and thereby discourage any redundant tokens to be generated via autoregressive generation. $\alpha=r_{+}-r_{0}.$ (2) Curriculum Prompt Schedule: The number of easy prompts $q_{\theta,G}$ grows in step with the progressive decline in remaining redundancy.
@@ 4 DeCS
Given the above analysis, we propose DeCS, which contains three main designs to achieve the highest efficacy-efficiency tradeoff. First, to ensure that redundant tokens are penalized deterministically, we train a small module that precisely identifies necessary reasoning prefix (NRP) components for correct trajectories (§4.1). After that, we design a decoupled token-level reward and differentiate the reward scale for necessary and redundant tokens, to ensure enhanced efficiency without compromising performance (§4.2). Based on the conception of NRP, we propose to prevent aggressive penalization on high-entropy tokens following NRP by refactoring the data distribution of a batch according to the current levels of redundancy incrementally (§4.3). Fig. 2 illustrates the overall algorithm.
@@ 4.1 Detection of NRP
Here, $\bigoplus$ denotes the concatenation of reasoning chunks, and the $c^{*}$-th reasoning chunk is the first to entail the correct answer $y^{*}$. The training details are illustrated in Appendix G.6.
@@ 4.2 Decoupled Reward Assignment
$r_{i,j}=\begin{cases}r_{+}\cdot\bm{1}_{{\bm{o}}_{i}\text{ is correct}}\quad&j\leq K_{o_{i}}^{*}\vee o_{j}\notin o_{\rm think}\vee o_{i,j}=\emptyset\\ (r_{0}-\frac{(r_{+}-r_{0})L_{i}}{L_{\max}})\cdot\bm{1}_{{\bm{o}}_{i}\text{ is correct}}\quad&j>K_{o_{i}}^{*}\land o_{j}\in o_{\rm think}\end{cases}$ (9)
@@
Appendix C in detail explains the functionality of Eq. 9 for penalizing any leading redundant tokens.
@@ 4.3 Curriculum Prompt Schedule
where $\kappa_{m}^{0}$ is the ratio of $q_{\theta,G}$ among the current sampled batch and $\beta$ is a hyperparameter to control the learning progress. As trajectories with zero advantages would not provide any learning signal, we follow Yu et al. (2025) to filter prompts whose $G$ rollouts are all incorrect and fill the batch by over-sampling. This curriculum strategy, designed to be bounded and monotonic, enables smooth adjustment in response to the observed NRP ratio, which aligns with the principle of curriculum learning (Bengio et al., 2009). By setting a moderate value $\beta$ with grid search (see Appendix H.1), DeCS can satisfy the condition elucidated in Theorem 1 to maintain unbiased learning of high-entropy tokens throughout the whole training process. This yields stable convergence with no observed training instability or performance degradation, which is reflected in Fig. 6 and 7.
@@ 5.1 Experiment Setups
We use MATH500 (Lightman et al., 2023), AMC23 (AI-MO, 2024), OlympiadBench (He et al., 2024), AIME2024 (Mathematical Association of America, 2025a) and AIME2025 (Mathematical Association of America, 2025b) as in-domain testbeds, GPQA-Diamond (GPQA-D; Rein et al. (2024)) and LiveCodeBench-v6 (LCB; Jain et al. (2025)) as held-out testbeds, covering math, coding, and science tasks with diverse complexity. We choose ThinkPrune (Hou et al., 2025), TLMRE (Arora and Zanette, 2025), AdaptThink (Zhang et al., 2025b), LC-R1 (Cheng et al., 2025) as baselines, and also include GRPO to serve as a performance reference. For fair comparison, we set the temperature as 0.6, top_p as 0.95, and use a maximum token limit of 16384 suggested by Guo et al. (2025). We conduct 128 rollouts for AIME2024, AIME2025, and AMC23, 16 rollouts for OlympiadBench, MATH500 and GPQA-D, and 10 rollouts for LCB to compute pass@1. We also compute the Average Efficiency Score (AES; Luo et al. (2025a)) for a comprehensive assessment of efficiency and efficacy. The details of both metrics are presented in Appendix G.4.
@@ 5.3 Ablation Study
5.4 Backbone Generalization
In this section, we generalize DeCS to Qwen3 backbone model, where we apply DeCS to Qwen3-4B (Yang et al., 2025) with the same training hyperparameters introduced in §5.1. Results in Table 2 demonstrates that DeCS successfully extends to Qwen3-4B, with 54.80% reduction of reasoning tokens and 1.32 pass@1 improvement on average. This strongly implies that DeCS is backbone-robust, and remains effective on a stronger base model.
Table 2: Generalization to the Qwen3-4B model. DeCS still achieves 0.61 AES score, with 54.80% reduction to overthinking and 1.32 pass@1 improvement.
Model AIME2024 AIME2025 AMC23 MATH500 OlympiadB GPQA-D LCB Average
Qwen3-4B 64.82 11611 56.30 12870 91.60 7478 93.74 4839 71.07 9144 39.11 8072 62.17 9713 68.40 9104 0.00
DeCS 65.38 5431 56.96 5758 93.59 2864 93.78 1648 74.09 3646 41.00 3260 63.27 6196 69.72 4115 0.61
@@ 7 Conclusion
Acknowledgments
We thank the anonymous reviewers for their insightful comments and suggestions. This work was supported by the National Key R&D Program of China (No. 2022ZD0162101), National Natural Science Foundation of China (No. 62576209) and STCSM (No. 2025SHZDZX025G05).
@@ 8 Reproducibility Statement
    Datasets: The training set we used is described in §5.1 and evaluation sets we used are described in Appendix G.2.
@@ A.3 Proof of Theorem 2
Table 3: Ablation study with two major components of DeCS on the DS-1.5B base model. “CS” denotes adaptive data sampling and “DR” denotes the decoupled reward mechanism.
@@ Appendix B Related Work
As is introduced in the main text, length reward is one of the most influential strategies to improve efficiency for large reasoning models (LRM). Su and Cardie (2025) proposes to determine the reward scale coefficient $\gamma$ defined in Eq. 4 by comparing the average accuracy of the current step with that of the reference model. Aggarwal and Welleck (2025) proposes L1, which optimizes the LRM to generate a correct reasoning path within a certain context window. Lyu et al. (2025) also proposes to use length penalties to encourage the model to answer problems with different complexity with adaptive token limits. Cheng et al. (2025) uses a specialized model to separate the necessary reasoning part and the redundant tokens, and reinforce the policy to output the response with the shortest NRP length. However, although they have made great efforts to encourage correct yet short reasoning paths, they do not consider the internal logit dynamics when applying sequence-level length rewards with token-level optimization objectives, and thus suffer from performance degradation.
@@
Table 4: Concepts of involved mathematical notations, symbols and abbreviations.
Symbol Description
$\pi_{\theta}$ The policy model parameterized by $\theta$.
$q$ A given input prompt or question.
$\bm{o}$} A reasoning trace (output sequence) generated by the policy $\pi_{\theta}$.
${\bm{o}}_{<j}$ The sub-sequence of the output $o$ up to (but not including) the $j$-th token.
$o_{j}$ The $j$-th token in the output sequence $o$.
$r$ The reward signal received from the verifier.
$\nabla J(\theta)$ The gradient of the policy objective function with respect to parameters $\theta$.
$A({\bm{o}}_{<j},j)$ The advantage function at step $j$ given the state ${\bm{o}}_{<j}$, indicating how much better the action is compared to average.
NRP Necessary Reasoning Prefix, the minimal set of prefix tokens of a reasoning trace required to reach a correct answer. Formal definition in Definition 1.
PNRP Proportion of the NRP tokens among a complete reasoning trace.
$L_{i},\left|{\bm{o}}_{i}\right|$ The total length of the $i$-th rollout (total number of tokens generated).
$\kappa_{m}$ The proportion of easy prompts among a batch at iteration $m$
$\beta$ Coefficient for increasing the $\kappa$ value for each iteration.
$r_{+}$ Maximum reward assigned to necessary reasoning tokens.
$r_{0}$ Base reward value.
AES Average Efficiency Score, a metric combining pass@1 score and token cost.
$\oplus$ Concatenation operator for reasoning chunks.
$c^{*}$ The index of the first reasoning chunk that entails the correct answer.
$K_{c^{*}}^{i}$ The cumulative token count up to the end of the $c^{*}$-th chunk in the $i$-th rollout.
Appendix C Detailed Analysis of Decoupled Reward Design
While the decoupled reward formulation in Eq. 9 does not explicitly differentiate between leading redundant tokens (i.e., the first token immediately following the Necessary Reasoning Prefix, NRP) and other redundant tokens, the combination of this design with the group-relative advantage mechanism in DECS ensures that all redundant tokens consistently receive negative advantages during training. This property arises from the interplay between the reward structure and the relative advantage computation:
  1. 1.
    Reward Structure: For any correct rollout $o_{i}$, tokens within the NRP receive a fixed high reward $r_{+}=1.1$. Tokens after the NRP (redundant tokens) receive a length-scaled reward:
    $r_{i,j}=r_{0}-\frac{(r_{+}-r_{0})L_{i}}{L_{\max}},\quad\text{where }L_{i}=|o_{i}|.$
  2. 2.
    Group-Relative Advantage Mechanism: A token receives positive advantage only if its reward exceeds the average reward at that position across the group of $G$ rollouts.
Consider the leading redundant token in a sequence:
Case 1: The current sequence has the longest NRP in the group.
Its reward computed by Eq. 9 is generally below-average the average reward among the token group. In this case, it receives negative advantages. This could be empirically verified across different training steps and model architectures. To be more specific, we evaluated rollouts from the DS-1.5B/DS-7B model at training steps [1, 60, 120, 180, 240] and counted how many leading redundant tokens have an above-average reward computed by Eq. 9 in Table C. Results demonstrate that across all batches and positions, no instance was found where a leading redundant token received a positive advantage.
Case 2: At least one sequence in the group has an equally long or longer NRP.
Then, there exists at least one token in that sequence receiving the full reward $r_{+}=1.1$. Meanwhile, under the condition where a longer sequence has longer NRP length, we could sort the 16 sequence lengths in descending order and assume that for a sequence length $L_{k}$ which is the $k$-th largest among the remaining 15 sequences. In this situation, there would be at least $k$ rewards that equal to $1.1$ and the remaining rewards are all less than or equal to $r_{k}$. Therefore,
$r_{k}-\mu=1.0-0.1\times L_{k}/L_{\max}-\frac{1}{16}(1.1\times k+\sum_{i=1}^{16-k}1.0-0.1\times L_{i}/L_{\max})$
where $\forall i,L_{i}\leq L_{k}$ by above conditions. Simplifying the right term, we could obtain:
$r_{k}-\mu$ $\leq 1-0.1\frac{L_{k}}{L_{\max}}-1-0.1\frac{k-(16-k)\frac{L_{k}}{L_{\max}}}{16}$
$=-0.1\frac{k}{16}(1+\frac{L_{k}}{L_{\max}})$
$<0$
This holds for any leading redundant tokens for any sequence that does not have longest NRP length, which also represents a negative advantage value.
In both cases, the reward for the leading redundant token is strictly below the group average, resulting in a negative advantage. This guarantees it will be penalized by the policy gradient update. Thus, the decoupled reward design, in conjunction with the advantage estimator of GRPO, inherently penalizes leading redundant tokens, thereby penalizing all redundancies by autoregressiveness.
Table 5: Comparison of erronously rewarded redundant tokens using Eq. 9 on DS-1.5B/DS-7B.
Step 1 60 120 180 240
DS-1.5B 0 0 0 0 N/A
DS-7B 0 0 0 0 0
Table 6: Human evaluation results on the classification accuracy of math-specialized NRP detector on other domains, including science and coding.
Dataset Human1 Human2 Human3
GPQA-D 99.18 98.96 99.23
LiveCodeBench 97.88 97.63 97.49
@@ Appendix C Detailed Analysis of Decoupled Reward Design
Appendix D Correlation to Relative Overgeneralization in Multi-Agent RL
In this section, we discuss the parallels between the “erroneous penalization” issue presented in this paper and concepts like relative overgeneralization in Multi-Agent RL (MARL). The core problem, the misalignment between global reward signals and local token-level updates, resonates with broader challenges in the RL literature.
  • The Parallel: Relative overgeneralization in MARL occurs when agents learn suboptimal joint behaviors due to misleading credit assignment from shared rewards. Analogously, in our single-agent sequence setting, the coarse-grained, sequence-level length penalty acts as a blunt signal. This signal fails to accurately attribute cost to specific redundant tokens, leading to the erroneous suppression of useful, high-entropy tokens. This is essentially a form of token-level overgeneralization.
  • The Solution: This parallel highlights a fundamental challenge in policy gradient methods: the need for fine-grained, temporally precise feedback to avoid spurious credit assignment. Our proposed decoupled reward mechanism (Eq. 9) can be viewed as an instance of structured credit assignment. By explicitly disentangling necessary and redundant reasoning steps, we ensure that only truly redundant tokens are penalized, effectively mitigating this token-level overgeneralization.
Appendix E Limitation & Future Work